OpenStudy (anonymous):

lim [(4^n + 5^n)^1/n] = ? n->infinity

OpenStudy (jtvatsim):

I'm thinking logarithmic might be the way to go here... not sure yet.

OpenStudy (anonymous):

I took the logarithm. After that should I proceed with L Hospital rule? Although I did it but didn't arrive to the result. Please help

OpenStudy (kc_kennylau):

What did you get when using l'hopital?

OpenStudy (kc_kennylau):

=lim[n->infty] e^{[ln(4^n+5^n)]/n} =e^{lim[n->infty][ln(4^n+5^n)]/n} =e^{lim[n->infty](4^nln4+5^nln5)/(4^n+5^n)}

OpenStudy (anonymous):

yes, I got this exctly. But how to proceed further?

OpenStudy (kc_kennylau):

Well.......

OpenStudy (kc_kennylau):

Any people here know?

OpenStudy (jtvatsim):

yikes... that looks nasty :P

OpenStudy (jtvatsim):

lemme see, I think I might have something...

OpenStudy (jtvatsim):

\[\lim_{n\rightarrow \infty} (4^n + 5^n)^{1/n} = \lim_{n\rightarrow \infty}(5^n[(\frac{4}{5})^n + 1^n])^{1/n} = \lim_{n \rightarrow \infty}5[(\frac{4}{5})^n + 1]^{1/n} = 5 \lim_{n\rightarrow \infty}[(\frac{4}{5})^n + 1]^{1/n}\]

OpenStudy (kc_kennylau):

Is that a fail

OpenStudy (jtvatsim):

so now, we use ln on the remaining limit portion

OpenStudy (jtvatsim):

\[\lim_{n \rightarrow \infty} \frac{1}{n}\ln(1 + [4/5]^n)=0\] because ln(1) = 0.

OpenStudy (jtvatsim):

This gives the final answer of \[5 e^{limit thing} = 5e^0 = 5\]

OpenStudy (anonymous):

OK, got that clearly. Thank you.

OpenStudy (jtvatsim):

no problem, that was tricky :P

OpenStudy (kc_kennylau):

Wow